OpenStudy (anonymous):

Find two geometric means between 7 and 875.

OpenStudy (anonymous):

you have to write it like this 7/x = 875/x wherein X is the geometric mean To find X use cross multiplication. It's just simple algebra

OpenStudy (anonymous):

oh sorry the right equation will be 7/x= x/875

OpenStudy (anonymous):

those are the geometric means?

OpenStudy (anonymous):

The value you will get for X is the geometric means

OpenStudy (anonymous):

im still not getting it

OpenStudy (amistre64):

essentially youre creating a geometric sequence

OpenStudy (amistre64):

a1 = 7 and a4 = 875 find a2 and a3

OpenStudy (amistre64):

an = a1 r^(n-1) a4 = 7 r^3 = 875 r = (875/7)^(1/3) a2 = 7 * (875/7)^(1/3) a3 = 7 * (875/7)^(2/3) if i see it right

OpenStudy (amistre64):

otherwise you can try to solve:\[\frac{7}{a_2}=\frac{a_2}{a_3}=\frac{a_3}{875}\]